# Spectral-Operator Calculus (Part A): Trace-Form Evaluators and Spectral Growth Taxonomy

### v1

@@ Abstract.
We develop a spectral-operator calculus on separable Hilbert space that treats self-adjoint operators and their bounded spectral transforms as the basic objects. On a class of such “spectral geometries” we introduce abstract evaluators, required to satisfy natural invariance, locality, extensivity, and dominated-convergence continuity conditions. Our first main result is a trace-form representation theorem: on the natural trace-class envelope every such evaluator is given by the trace of a single nondecreasing profile applied through the functional calculus. Thus, once that profile and a scalar normalization are fixed, all admissible scalar values are determined by the underlying spectra, yielding a rigidity principle for evaluators in this setting. The second main result is a spectral growth taxonomy: we classify self-adjoint operators by counting-function asymptotics and show that the polynomial growth class is stable under the basic constructions of the calculus. Together these results provide an arithmetic-neutral analytic backbone for later parts of the series and for applications to concrete spectral models.
@@
Series note. This article is the first part of a planned sequence developing a spectral-operator calculus for self-adjoint operators on separable Hilbert space. Part A establishes the evaluator axioms, proves a trace-form representation theorem on a natural trace-class envelope, and introduces a spectral growth taxonomy and continuity framework for scalar evaluations built from spectra. Later parts use this analytic backbone to study admissible kernels, trace cutoffs, analytic continuation, and boundary reconstruction, but the present paper is logically self-contained and can be read independently of any later developments in the series.
@@ 1.1. Main results and contributions
Our first main result is Theorem~3.1, a trace-form representation theorem for evaluators. Under the standing hypotheses on $\mathcal{C}$ and the evaluator axioms, there exist a Borel-measurable nondecreasing function $h:\mathbb{R}\to\mathbb{R}$ and a normalization constant $c_{\mathcal{E}}\geq 0$ such that $\mathcal{E}(f(D))$ is given by the trace of $h(f(D))$ whenever this operator is trace class. In particular, if we restrict $\mathcal{E}$ to the trace-class envelope
@@
The second main result is a spectral growth taxonomy for self-adjoint operators. We work with counting functions $N_{D}(\Lambda)$ and related spectral distributions, and use their asymptotic behaviour as $\Lambda\to\infty$ to organize operators into growth classes. In particular, we single out a polynomial growth regime that captures the standard elliptic and Laplace-type models and show in Theorem~3.5 that this class is stable under the basic constructions of the calculus, including monotone regularly varying functional calculus, finite orthogonal sums, and tensor products. This growth framework is then used in examples and in the continuity theory for evaluators in Section 5.
@@ 1.2. Relation to prior work
There is a substantial literature on trace ideals, unitarily invariant norms, and operator inequalities, with classical references including Reed–Simon, Simon’s monograph on trace ideals, and Bhatia’s text on matrix analysis. In that setting one typically starts from a given trace and studies the properties of associated norms and functionals on operator ideals. The perspective adopted here is different: we start from an abstract evaluator $\mathcal{E}$ defined on a class $\mathcal{C}$ of spectral geometries, subject only to unitary invariance, projector-locality, dominated continuity, and a growth bound, and obtain a universal trace-form representation for such evaluators on a trace-class envelope. In this sense Theorem~3.1 can be viewed as a representation theorem for evaluator functionals, rather than as a study of a fixed underlying trace.
@@
On the functional-calculus side, our axiomatic stance is parallel to the general theory of unbounded functional calculi and spectral integrals, as developed for example in Haase’s work on sectorial operators and general functional calculus frameworks. In such settings one studies large classes of operators admitting a rich functional calculus and then constructs spectral integrals with respect to operator-valued measures. The present paper is complementary: we work within the usual spectral theorem for self-adjoint operators, but impose additional structural conditions on scalar-valued evaluators acting on bounded transforms. Theorem~3.1 then shows that, under these axioms, such evaluators are necessarily of trace form on $\mathcal{C}_{1}$.
On the growth side, the taxonomy introduced here is influenced by the language of regular variation and spectral asymptotics. Counting-function techniques and regularly varying functions are standard tools in the analysis of eigenvalue distributions, as in the work of Bingham–Goldie–Teugels and in more recent treatments of spectral densities and hypertraces in noncommutative geometry, as well as in explicit counting-function studies for fractal sprays and related models. Our contribution is to formulate a growth classification directly in terms of counting functions and to prove stability of the polynomial class under the structural operations required by the evaluator calculus, such as functional calculus, orthogonal sums, and tensor products, without committing to a specific geometric model.
This paper should therefore be read as an operator-theoretic calculus for evaluators and spectral growth, rather than as a contribution to the fine structure of trace ideals or spectral asymptotics. It does not aim to optimize asymptotic constants or to exhaust the classical theory; instead, it isolates a structural setting in which evaluators, spectral geometries, and growth conditions fit together coherently, and in which the main representation and continuity results can be applied across different models without further analytic reconstruction. The results of Part A are intended to be reusable: they provide an analytic backbone for later work on kernels, cutoff schemes, and flow constructions on spectral data, while the present part can be read independently of any later developments in the series.
@@ 2.1. Spectral measures and functional calculus
in the sense of strong operator convergence on the domain of $D$; see Reed-Simon [1, Theorem VIII.3]. For any bounded Borel function $f:\mathbb{R}\to\mathbb{C}$, the bounded operator
@@
When $D$ has pure-point spectrum with finite multiplicities, we write $\{\lambda_{n}\}_{n\geq 1}$ for its eigenvalues (repeated according to multiplicity) and choose an orthonormal basis of eigenvectors. Then $D$ is unitarily equivalent to a diagonal operator on $\ell^{2}$, and for bounded Borel $f$,
$f(D)\sim\operatorname{diag}\big(f(\lambda_{1}),f(\lambda_{2}),\ldots\big).$
@@ 2.3. Counting functions and growth
When $D$ has discrete spectrum with finite multiplicities accumulating only at $\pm\infty$, its eigenvalues can be numbered as $\{\lambda_{n}\}_{n\geq 1}$, and the associated counting function is
$N_{D}(\Lambda):=\#\{n\geq 1:|\lambda_{n}|\leq\Lambda\},\qquad\Lambda\geq 0,$
@@ 2.4. Spectral preorder and majorization
In this paper we use the spectral preorder $\preceq$ only for operators with discrete spectra and finite multiplicities, as in Definition 2.4. All subsequent appearances of $\preceq$ refer to that setting. Extensions of spectral majorization to more general self-adjoint operators via strong resolvent limits are well developed in the literature (see, for example, Bhatia [5, Chapter III]), but we do not rely on any such extension here, and no properties of $\preceq$ beyond the discrete-spectrum case are needed for the results of Sections 3-5.
@@ 2.5. Evaluators and structural notation
assigning a nonnegative extended real number to each bounded spectral geometry in $\mathcal{C}$. The structural axioms imposed on evaluators – unitary invariance, extensivity on orthogonal sums, projector-locality, dominated continuity, and a normalization/growth bound – will be stated and fixed in Section 3.1 as Axioms A1–A5. Until then, $\mathcal{E}$ is only used as notation for a generic scalar-valued functional on $\mathcal{C}$.
@@ 3.1. Standing assumptions and axioms
In this section we fix the structural framework for the evaluator calculus used throughout the paper. All later results in Part A are formulated under these standing assumptions.
@@
These axioms fix the structural setting for the trace-form representation theorem and the growth taxonomy developed in the rest of the paper. In particular, they encode the requirements that any lawful evaluator must satisfy in order to behave well under unitary conjugation, orthogonal decomposition, spectral partitions, and dominated limits. Under Axioms A1–A5, Theorem 3.1 will show that on $\mathcal{C}_{1}$ the evaluator $\mathcal{E}$ is necessarily of trace form with a single Borel nondecreasing profile and a global normalization constant, and Section 3.3 will introduce the growth classes compatible with this framework.
@@ 3. $D$ has logarithmic (or slower) growth if $N_{D}(\Lambda)\to\infty$ and either
We write $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\log}$ for the families of self-adjoint operators (and the spectral geometries they generate) with polynomial, stretched-exponential, and logarithmic-or-slower growth, respectively. These families are disjoint: no operator can belong to more than one class; see Section 4.2 and Appendix B for a justification based on regular variation.
@@
Theorem 3.5 shows that, in particular, the polynomial growth class $\mathcal{G}_{\mathrm{poly}}$ is stable under the basic constructions of the spectral calculus that we use later. This stability under monotone functional calculus, orthogonal sums, and tensor products underlies the continuity and evaluator examples in Sections 4-5.
@@ 4. Proofs of the main results
In this section we prove the representation and growth theorems stated in Section 3. Section 4.1 outlines the proof of Theorem~3.1 (Trace-form representation) and reduces the measure-theoretic details to Appendix A. Section 4.2 proves Theorem~3.5 (Closure of spectral growth classes), using the counting-function and regular-variation estimates collected in Appendix B.
4.1. Proof of Theorem~3.1 (Trace-form representation)
@@
By positivity, extensivity on orthogonal sums, projector-locality, and dominated continuity (applied to increasing sequences of projections dominated by a fixed spectral projection), the set function $B\mapsto\mu_{D}(B)$ extends from finite spectral partitions to a countably additive, finite measure on the $\sigma$-algebra generated by those spectral subsets, and $\mu_{D}(B)=0$ whenever $\nu_{D}(B)=0$. In particular, $\mu_{D}\ll\nu_{D}$; see Appendix A.2-A.3 for details.
@@
On the trace-class ideal $\mathcal{S}_{1}(H)$, the standard representation theorem for normal functionals on $B(H)$ (see Simon [6, Chapter 1]) yields a unique positive trace-class operator $T_{D}$ such that
@@
for some function $h:\mathbb{R}\to\mathbb{R}$ determined by the values of $\mathcal{E}$ on rank-one spectral blocks. Using projector-locality and dominated continuity, Appendix A.2-A.4 extend this construction to positive trace-class operators $X\in\mathcal{C}_{1}^{+}$ by approximating $X$ from below with finite-rank truncations and passing to the limit under domination.
@@
for every $X\in\mathcal{C}_{1}^{+}$. Normalization and uniqueness (Appendix A.5-A.6) then show that there exists a constant $c_{\mathcal{E}}>0$, independent of $D$ and $f$, such that
@@
4.2. Proof of Theorem~3.5 (Closure of spectral growth classes)
We keep the notation of Section 3.3. Throughout, $D$ and $E$ are self-adjoint operators with discrete spectra and counting functions $N_{D}$ and $N_{E}$ satisfying Convention~2.2. The counting-function reparametrization and tensor-product estimates we need are collected in Appendix B. In Section 4.2.1 we restate the monotone reparametrization lemma from Appendix B.2 as Lemma~4.1. In Section 4.2.2 we prove the orthogonal-sum and tensor-product bounds; the polynomial tensor-product estimate is an immediate consequence of the convolution bounds in Appendix B.3.
@@ 4.2.1. Monotone functional calculus
We first record the counting-function identity underlying part (1) of Theorem~3.5.
@@ 4.2.3. Conclusion of the proof
The disjointness of the growth classes $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\log}$ follows from the incompatibility of their asymptotic forms. For example, if $N_{D}(\Lambda)\sim C\,\Lambda^{d}$ with $C,d>0$ and also $N_{D}(\Lambda)\sim\exp(c\,\Lambda^{\alpha})$ with $c>0$ and $0<\alpha<1$, then comparing logarithms as $\Lambda\to\infty$ yields a contradiction. More generally, the classification of regularly varying functions shows that polynomial, stretched-exponential, and logarithmic-or-slower growth regimes are mutually exclusive; see Appendix B and Bingham-Goldie-Teugels [5, Theorem 1.3.1 and Section 1.4].

### v2

# Spectral-Operator Calculus I: Trace-Form Evaluators and Spectral Growth in the Self-Adjoint Setting

@@ Abstract.
We develop Spectral-Operator Calculus (SOC), an axiomatic calculus for scalar evaluation of operator-generated spectral observables. This paper (SOC-I) treats the self-adjoint setting, where observables are bounded Borel transforms and locality is enforced via additivity across spectral partitions. Under unitary invariance, extensivity on orthogonal sums, projector-locality, and a dominated-convergence continuity condition, we prove a rigidity theorem on a natural trace-class envelope: every admissible evaluator agrees with a weighted trace of a single Borel nondecreasing profile applied through the functional calculus. We then introduce a spectral growth taxonomy based on eigenvalue counting asymptotics and show that the polynomial growth regime is stable under the basic constructions of the calculus. These results supply an arithmetic-neutral analytic backbone for subsequent SOC parts and for applications to concrete spectral models. A companion part treats the sectorial/holomorphic setting, where locality is formulated on log-scale via scale-band decompositions and positive kernels rather than spectral projections.
@@
Series note. The SOC series develops an evaluator calculus for spectral observables generated by functional calculi. The common spine is an axiomatics for scalar evaluators (unitary invariance, extensivity, localization under coarse-graining, dominated-continuity, and normalization), together with rigidity and stability consequences that are designed to persist under admissible limiting schemes.
SOC-I develops the self-adjoint/PVM-localized model case. It fixes the evaluator axioms, proves a trace-form rigidity theorem on a natural trace-class envelope, and introduces a spectral growth taxonomy and continuity framework for scalar evaluations built from spectra. A companion part develops the sectorial/holomorphic model case (Haase-type functional calculus; see Haase [1]), where the localization gate is implemented on log-scale by scale-band decompositions and kernel positivity rather than by spectral projections.
Later parts build on these model cases to study admissible kernels and cutoff schemes, analytic continuation procedures, and boundary and reconstruction questions.
@@ 1. Introduction
At the level of structure, SOC is organized around three interacting gates: (i) a localization rule that specifies how evaluator values decompose under coarse-graining; (ii) a dominated-continuity rule that legalizes passage to limits inside the evaluator; and (iii) a normalization/nontriviality rule that excludes degenerate assignments. The operator class and functional calculus supply the observable family and the available coarse-grainings. In the self-adjoint setting of this part, coarse-graining is spectral: we partition the spectrum and demand additivity across the corresponding spectral projections. In the sectorial setting, spectral projections are unavailable in general; SOC instead uses scale-band decompositions $t\mapsto\psi(tA)$ and controls interference through a positive kernel on log-scale. This paper develops the PVM-localized model case and proves that the gates force trace-form rigidity on the trace-class envelope.
@@ 1.1. Main results and contributions
Our first main result is Theorem 3.1, a trace-form representation theorem for evaluators. Under the standing hypotheses on $\mathcal{C}$ and the evaluator axioms, there exist a Borel-measurable nondecreasing function $h:\mathbb{R}\to\mathbb{R}$ and a normalization constant $c_{\mathcal{E}}\geq 0$ such that $\mathcal{E}(f(D))$ is given by the trace of $h(f(D))$ whenever this operator is trace class. In particular, if we restrict $\mathcal{E}$ to the trace-class envelope
@@
The second main result is a spectral growth taxonomy for self-adjoint operators. We work with counting functions $N_{D}(\Lambda)$ and related spectral distributions, and use their asymptotic behaviour as $\Lambda\to\infty$ to organize operators into growth classes. In particular, we single out a polynomial growth regime that captures the standard elliptic and Laplace-type models and show in Theorem 3.5 that this class is stable under the basic constructions of the calculus, including monotone regularly varying functional calculus, finite orthogonal sums, and tensor products. This growth framework is then used in examples and in the continuity theory for evaluators in Section 5.
@@ 1.2. Relation to prior work
There is a substantial literature on trace ideals, unitarily invariant norms, and operator inequalities, with classical references including Reed-Simon [2], Simon’s monograph on trace ideals [3], and Bhatia’s text on matrix analysis [4]. In that setting one typically starts from a given trace and studies the properties of associated norms and functionals on operator ideals. The perspective adopted here is different: we start from an abstract evaluator $\mathcal{E}$ defined on a class $\mathcal{C}$ of spectral geometries, subject only to unitary invariance, projector-locality, dominated continuity, and a growth bound, and obtain a universal trace-form representation for such evaluators on a trace-class envelope. In this sense Theorem 3.1 can be viewed as a representation theorem for evaluator functionals, rather than as a study of a fixed underlying trace.
@@
On the functional-calculus side, SOC is modular: it takes as input an operator class equipped with a functional calculus and a compatible notion of coarse-graining, and it studies which scalar evaluators are forced by invariance, locality, and dominated-limit legality. SOC-I instantiates this program in the self-adjoint/Borel setting, where the spectral theorem supplies both the calculus and the canonical coarse-grainings (spectral partitions via the PVM). Haase’s framework for sectorial operators [1] supplies the corresponding input for a second SOC model case based on holomorphic functional calculus, where PVM locality is replaced by scale-band decompositions and kernel positivity on log-scale. The novelty in SOC is not the existence of these calculi, but the rigidity and stability constraints on evaluators that follow once a localization structure is fixed.
On the growth side, the taxonomy introduced here is influenced by the language of regular variation and spectral asymptotics. Counting-function techniques and regularly varying functions are standard tools in the analysis of eigenvalue distributions, as in the work of Bingham-Goldie-Teugels [5] and in more recent treatments of spectral densities and hypertraces in noncommutative geometry [6], as well as in explicit counting-function studies for fractal sprays and related models [7]. Our contribution is to formulate a growth classification directly in terms of counting functions and to prove stability of the polynomial class under the structural operations required by the evaluator calculus, such as functional calculus, orthogonal sums, and tensor products, without committing to a specific geometric model.
This paper should therefore be read as an operator-theoretic calculus for evaluators and spectral growth, rather than as a contribution to the fine structure of trace ideals or spectral asymptotics. It does not aim to optimize asymptotic constants or to exhaust the classical theory; instead, it isolates a structural setting in which evaluators, spectral geometries, and growth conditions fit together coherently, and in which the main representation and continuity results can be applied across different models without further analytic reconstruction. The results of SOC-I are intended to be reusable: they provide an analytic backbone for later work on kernels, cutoff schemes, and flow constructions on spectral data, while the present part can be read independently of any later developments in the series.
@@ 2.1. Spectral measures and functional calculus
in the sense of strong operator convergence on the domain of $D$; see Reed-Simon [2]*{Theorem VIII.3}. For any bounded Borel function $f:\mathbb{R}\to\mathbb{C}$, the bounded operator
@@
When $D$ has pure-point spectrum with finite multiplicities, we write $\{\lambda_{n}(D)\}_{n\geq 1}$ for its eigenvalues (repeated according to multiplicity) and choose an orthonormal basis of eigenvectors. Then $D$ is unitarily equivalent to a diagonal operator on $\ell^{2}$, and for bounded Borel $f$,
$f(D)\sim\operatorname{diag}\big(f(\lambda_{1}(D)),f(\lambda_{2}(D)),\ldots\big).$
@@ 2.3. Counting functions and growth
When $D$ has discrete spectrum with finite multiplicities accumulating only at $\pm\infty$, its eigenvalues can be numbered as $\{\lambda_{n}(D)\}_{n\geq 1}$, and the associated counting function is
@@ 2.4. Spectral preorder and majorization
In this paper we use the spectral preorder $\preceq$ only for operators with discrete spectra and finite multiplicities, as in Definition 2.4. All subsequent appearances of $\preceq$ refer to that setting. Extensions of spectral majorization to more general self-adjoint operators via strong resolvent limits are well developed in the literature (see, for example, Bhatia [4]*Chapter III), but we do not rely on any such extension here, and no properties of $\preceq$ beyond the discrete-spectrum case are needed for the results of Section 3, Section 4, and Section 5.
@@ 2.5. Evaluators and structural notation
assigning a nonnegative extended real number to each bounded spectral geometry in $\mathcal{C}$. The structural axioms imposed on evaluators - unitary invariance, extensivity on orthogonal sums, projector-locality, dominated continuity, and a normalization/growth bound - will be stated and fixed in Section 3.1 as Axioms A1-A5. Until then, $\mathcal{E}$ is only used as notation for a generic scalar-valued functional on $\mathcal{C}$.
@@ 3.1. Standing assumptions and axioms
In this section we fix the structural framework for the evaluator calculus used throughout the paper. All later results in SOC-I are formulated under these standing assumptions.
@@
These axioms fix the structural setting for the trace-form representation theorem and the growth taxonomy developed in the rest of the paper. In particular, they encode the requirements that any lawful evaluator must satisfy in order to behave well under unitary conjugation, orthogonal decomposition, spectral partitions, and dominated limits. In terms of the three-gate structure of Section 1: Axioms A1 and A3 implement the localization gate; Axiom A4 implements dominated-continuity; and Axiom A5 implements normalization, while Axiom A2 (extensivity) is a structural compatibility condition. Under Axioms A1-A5, Theorem 3.1 will show that on $\mathcal{C}_{1}$ the evaluator $\mathcal{E}$ is necessarily of trace form with a single Borel nondecreasing profile and a global normalization constant, and Section 3.3 will introduce the growth classes compatible with this framework.
@@ 3. $D$ has logarithmic (or slower) growth if $N_{D}(\Lambda)\to\infty$ and either
We write $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ for the families of self-adjoint operators (and the spectral geometries they generate) with polynomial, stretched-exponential, and logarithmic-or-slower growth, respectively. These families are disjoint: no operator can belong to more than one class; see Section 4.2 and Appendix B for a justification based on regular variation.
@@
Theorem 3.5 shows that, in particular, the polynomial growth class $\mathcal{G}_{\mathrm{poly}}$ is stable under the basic constructions of the spectral calculus that we use later. This stability under monotone functional calculus, orthogonal sums, and tensor products underlies the continuity and evaluator examples in Section 4 and Section 5.
@@ 4. Proofs of the main results
In this section we prove the representation and growth theorems stated in Section 3. Section 4.1 outlines the proof of Theorem 3.1, with the main construction deferred to Appendix A. Section 4.2 proves the closure of the growth families in Theorem 3.5, with supporting estimates in Appendix B.
4.1. Proof of Theorem 3.1 (Trace-form representation)
@@
By positivity, extensivity on orthogonal sums, projector-locality, and dominated continuity (applied to increasing sequences of projections dominated by a fixed spectral projection), the set function $B\mapsto\mu_{D}(B)$ extends from finite spectral partitions to a countably additive, finite measure on the $\sigma$-algebra generated by those spectral subsets, and $\mu_{D}(B)=0$ whenever $\nu_{D}(B)=0$. In particular, $\mu_{D}\ll\nu_{D}$; see Lemma A.7 and Remark A.8.
@@
On the trace-class ideal $\mathcal{S}_{1}(H)$, the standard representation theorem for normal functionals on $B(H)$ (see Simon [3]*Chapter 1) yields a unique positive trace-class operator $T_{D}$ such that
@@
for some function $h:\mathbb{R}\to\mathbb{R}$ determined by the values of $\mathcal{E}$ on rank-one spectral blocks. Using projector-locality and dominated continuity, Lemma A.2, Lemma A.3, and Lemma A.4 extend this construction to positive trace-class operators $X\in\mathcal{C}_{1}^{+}$ by approximating $X$ from below with finite-rank truncations and passing to the limit under domination.
@@
for every $X\in\mathcal{C}_{1}^{+}$. Normalization and uniqueness (see Lemma A.5) then show that there exists a constant $c_{\mathcal{E}}>0$, independent of $D$ and $f$, such that
@@
4.2. Proof of Theorem 3.5 (Closure of spectral growth classes)
We keep the notation of Section 3.3. Throughout, $D$ and $E$ are self-adjoint operators with discrete spectra and counting functions $N_{D}$ and $N_{E}$ satisfying Convention 2.2. The counting-function reparametrization and tensor-product estimates we need are collected in Appendix B. In Section 4.2.1 we restate the monotone reparametrization lemma Lemma B.5 as Lemma 4.1. In Section 4.2.2 we prove the orthogonal-sum and tensor-product bounds; the polynomial tensor-product estimate is an immediate consequence of Lemma B.6 and Corollary B.8.
@@ 4.2.1. Monotone functional calculus
We first record the counting-function identity underlying part (1) of Theorem 3.5.
@@ 4.2.3. Conclusion of the proof
The disjointness of the growth classes $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ follows from the incompatibility of their asymptotic forms. For example, if $N_{D}(\Lambda)\sim C\,\Lambda^{d}$ with $C,d>0$ and also $N_{D}(\Lambda)\sim\exp(c\,\Lambda^{\alpha})$ with $c>0$ and $0<\alpha<1$, then comparing logarithms as $\Lambda\to\infty$ yields a contradiction. More generally, the classification of regularly varying functions shows that polynomial, stretched-exponential, and logarithmic-or-slower growth regimes are mutually exclusive; see Appendix B and Bingham-Goldie-Teugels [5]*Theorem 1.3.1 and Section 1.4.